# DarwinNet: An Evolutionary Network Architecture for Agent-Driven Protocol Synthesis

### 27 Mar 2026 v1

@@ I Introduction
For decades, the classical edifice of computer science has been built upon the aesthetics of completeness and determinism. Nowhere is this more evident than in the “narrow waist” of the TCP/IP architecture. While this rigid design enabled the global expansion of the Internet, it has inadvertently led to a phenomenon known as protocol ossification [1, 9]. In modern heterogeneous environments—ranging from high-latency satellite constellations to massive-scale IoT deployments—the standardization cycle of traditional protocols is prohibitively long, often taking years to ratify minor adjustments to headers or congestion control logic. Furthermore, the traditional paradigm of “Code is Law” assumes that engineers can pre-emptively define every combination of input and state during the design phase. However, this reliance on mechanical completeness creates inherently brittle systems. When faced with undefined edge cases or novel environmental stressors, rule-based systems either crash or fall into unknown error states, lacking the flexibility to adapt without human intervention.
As we enter the era of Large Language Models (LLMs) and autonomous Agents, the fundamental nature of a network node is changing [27, 20]. Computing nodes are no longer merely passive executors of deterministic instructions; they have evolved into intelligent agents capable of probabilistic reasoning and intent understanding. In this context, legacy deterministic communication protocols have become a bottleneck, sacrificing efficiency for human readability—a constraint we term Carbon Hegemony [^1] Here Carbon Hegemony refers to the constraints imposed by human-centric cognitive structures on machine-to-machine communication, rather than environmental carbon emissions.. DarwinNet seeks to dismantle this by envisioning a formal parting of ways between computer protocols and human language, allowing machines to communicate via high-dimensional logic that transcends human cognitive structures. Crucially, this evolution does not imply the abandonment of existing network foundations. Instead, DarwinNet is positioned as a high-order augmentation for the 6G era, designed to handle hyper-heterogeneous edge scenarios where traditional manual standardization reaches its scalability limits, while still relying on the global reachability provided by the TCP/IP ”narrow waist” as its foundational anchor.
@@
To realize the transition from structural fragility to evolutionary anti-fragility depicted in Fig. 1, this paper introduces DarwinNet—a bio-inspired, agent-driven network architecture that enables protocols to be “grown” rather than “designed.” Unlike traditional networks that succumb to ossification under static standards, DarwinNet conceptualizes the network as a living ecosystem capable of autonomous adaptation [24]. Structurally, this is achieved through a tri-layered decoupling: a Physical/Mathematical Anchor (Layer 0) for immutable constraints, a Fluid Cortex (Layer 1) serving as the adaptive execution body, and a Darwin Cortex (Layer 2) acting as the intelligent decision-making brain. By shifting the paradigm from rigid header-matching to high-dimensional intent negotiation, DarwinNet allows nodes to synthesize and deploy specialized communication logic in real-time. This mechanism ensures that network anomalies are no longer fatal errors, but rather catalysts for evolution that drive the system toward higher efficiency and survival.
In summary, DarwinNet represents a fundamental reconstruction of the network interaction model, moving from rigid rules to organic adaptation [11]. The core contributions of this work are as follows: 1. Paradigm Shift from Design to Growth:We propose a self-adaptive network architecture that achieves anti-fragility. By allowing protocols to evolve as fluid expressions of node intents, the network can autonomously adapt to environmental mutations—such as unforeseen congestion or novel attack patterns—without human-issued patches. 2. Intent-to-Bytecode (I2B) Synthesis with Dual-Loop Mechanism:We design an adaptive framework that leverages LLMs to transform high-level business intents into restricted, executable WebAssembly (WASM) bytecode. This is governed by a dual-loop architecture: an outer loop for cognitive protocol synthesis and an inner loop for millisecond-level Hot Swaps within a zero-trust sandbox [18]. 3. Protocol Solidification Index (PSI):We define a novel metric to quantify the maturity of the evolving system. This index captures the process of “Protocol Liquefaction,” where high-level reasoning (System 2) gradually collapses into highly efficient, compiled execution paths (System 1) as common interaction patterns stabilize. 4. Reliability Verification Paradigm:We introduce a mathematical verification framework based on the Crow-AMSAA (Duane) model. By treating protocol mutations as a reliability growth process, we provide a formal method to ensure that autonomous evolution remains within safety boundaries while continuously reducing failure rates over time.
@@ II Related Work
Programmable Data Planes and eBPF. The advent of P4 (Programming Protocol-independent Packet Processors) and eBPF (extended Berkeley Packet Filter) has shifted the networking paradigm from fixed-function hardware to programmable targets [4]. These technologies allow developers to define custom packet-processing logic and execute it at near-native speeds within the kernel or on specialized switching chips [4, 18]. However, existing programmable data planes primarily function as execution environments that require human-defined logic at compile-time or deployment-time. While they provide the necessary “body” for flexible networking, they lack an autonomous “brain” capable of real-time reasoning and logic synthesis [19]. DarwinNet builds upon these execution primitives but elevates them by introducing an agent-driven layer that dynamically generates and swaps protocol logic during runtime.
@@
Intent-Based Networking (IBN). Traditional IBN systems aim to bridge the gap between high-level business goals and low-level network configurations [15]. Most current IBN implementations are limited to “parameter-level” adjustments, such as automatically tuning TCP window sizes or re-routing paths based on pre-defined policies [19]. DarwinNet pushes the boundaries of IBN into the domain of “logic-level” synthesis [21, 20]. Instead of mapping intents to a library of static, human-written scripts, DarwinNet utilizes a “fluid cortex” to synthesize entirely new protocol behaviors in the form of executable bytecode. This transition from “Intent-to-Policy” to I2B enables a degree of architectural agility that conventional IBN cannot achieve [15, 5].
@@
Emergent Communication in Multi-Agent Reinforcement Learning (MARL). Research in MARL has demonstrated that autonomous agents can “invent” their own communication protocols to solve collaborative tasks, often resulting in high-dimensional vector-based languages that are unintelligible to humans [16, 22, 7, 17]. While these emergent protocols achieve high efficiency, they are often “black boxes” embedded within neural weights, making them difficult to verify, port, or debug. While maintaining safety via WASM, DarwinNet acknowledges that human experts are no longer the primary coders of protocol syntax. This marks a transition of computer science from a pure engineering discipline to an experimental science, where humans act as gardeners of an evolving ecosystem rather than micro-managers of bit-level headers. By using WASM as the synthesis target [31, 21], DarwinNet ensures that evolved protocols are explicit, portable, and subject to formal safety constraints within a zero-trust sandbox [31].
@@
Mimic Defense and Endogenous Security. The theory of Mimic Defense, characterized by Dynamic Heterogeneous Redundancy (DHR), suggests that system security can be enhanced by introducing uncertainty and dynamic transformations to thwart attackers. DarwinNet inherently realizes this security philosophy through the mechanism of protocol mutation [2]. By constantly evolving the syntax and semantics of communication—rather than relying on a static, standardized “instruction manual” like TCP/IP—DarwinNet creates a Moving Target Defense (MTD) at the protocol level. This dynamic nature significantly increases the cost of reconnaissance and exploitation for attackers, as the network’s “attack surface” is in a state of constant, intelligent flux.
@@
Recursive InterNetwork Architecture (RINA). RINA represents a fundamental rethink of the OSI model, proposing a recursive structure that treats networking as a single type of layer (the Distributed IPC Facility) repeated at different scales [12, 10]. While RINA provides an elegant theoretical framework for cross-layer optimization and de-layering, it still assumes a relatively stable set of mechanisms and policies. DarwinNet shares RINA’s vision of breaking down rigid layer boundaries but focuses more on the “biological” growth of functions rather than the structural recursion of the architecture [30]. DarwinNet’s “liquefaction” process treats protocol functions as atomic components that can be organically reassembled and evolved based on environmental fitness rather than predefined recursive templates.
@@ III-A Dual-Process Theory in Network Intelligence
The evolution of a protocol in DarwinNet is essentially the process of “liquid” System 2 reasoning collapsing into “solid” System 1 reflexes to achieve maximum efficiency [29, 23, 3].
@@ III-B WebAssembly and Sandbox Isolation
To ensure that autonomously synthesized logic does not compromise node security, we utilize WASM as the fluid execution target. WASM provides a memory-safe, hardware-independent, and near-native performance environment [18]. Its capability for millisecond-level instantiation and hot-swapping allows DarwinNet to perform logic mutations without interrupting active data streams. Furthermore, the use of restricted WASM bytecode facilitates formal verification of synthesized logic, ensuring it adheres to the “Constitution” defined in Layer 0 [31].
@@ III-C Crow-AMSAA Reliability Growth Model
The evolutionary maturity of DarwinNet is modeled using the Crow-AMSAA (Duane) Power Law process. This model describes the reliability growth of a system undergoing iterative “discovery and repair” cycles [8]. The failure (or protocol mismatch) intensity function is defined as:
@@ IV The Proposed Framework
This tri-layered, dual-path design ensures that DarwinNet can decouple the “intelligence of the protocol” from the “mechanics of the transmission,” allowing nodes to evolve their communication language without disrupting the underlying physical connectivity.
@@ IV-A Layered Architecture: From Anchor to Fluid
DarwinNet is organized into three distinct layers, transitioning from “Physical/Mathematical Hardness” at the base to “Intelligent Liquidity” at the apex.
@@
Layer 0 represents the fundamental “laws of physics” of the network that remain constant throughout the evolutionary process. This layer ensures basic bit-level connectivity (e.g., the standard TCP/IP stack for physical reachability) and enforces mathematical truths, such as cryptographic verification and logical consistency ($1+1=2$). In this paradigm, we relinquish micro-control over the communication process (the Silicon-driven protocol) and retain only mathematical constraints over the results (the Consensus). Layer 0 ensures that mathematics remains the only universal language between carbon-based intent and silicon-based execution, serving as the final tether of human control over autonomous evolution [10].
This layer functions as the ”Mathematical Constitution” of the network. No matter how complex or ”liquid” the evolved logic in Layer 1 becomes, it is strictly forbidden from violating the physical and security predicates pre-defined by human architects in Layer 0. This ensures that the autonomy of the network is always bounded by a ”Human-in-the-loop” safety framework, preventing any unpredictable or runaway protocol behavior.
@@
Layer 1 is the execution environment of the network, characterized by its “liquid” state. Unlike traditional fixed-protocol implementations, the Fluid Cortex utilizes a high-performance, zero-trust sandbox (e.g., WebAssembly) to run protocol logic. This layer serves as the “system’s body” or the “Fast Path” (System 1), executing synthesized bytecode at near-native speeds. It supports millisecond-level hot-swapping of logic—such as switching from a JSON-based parser to a specialized binary delta-compression algorithm—without interrupting data flows.
@@
As illustrated in Fig. 3, the Fluid Cortex is architected as a bionic Dual-Loop feedback system, conceptualizing the network node as a self-evolving “liquid organism” [25, 30]. This design explicitly bifurcates the processing logic into two cognitive pathways:
  • The Polymorphic Runtime (System 1): Occupying the lower tier, this “body” serves as the Fast Path. It utilizes deformable, liquid-like containers to carry real-time data streams. Its primary objective is the pursuit of ultimate execution efficiency, operating on the evolved “reflexes” (bytecode) generated by the upper layer.
  • The Darwin Cortex (System 2): Occupying the upper tier, this “brain” serves as the Slow Path. It is responsible for high-level perception, identifying shifts in user intent and environmental anomalies through its “sensory nerves.”
The synergy between these two loops enables a sophisticated evolutionary reflex: when the sensory nerves trigger an evolution request, the Darwin Cortex retrieves historical experiences from a Gene Pool and synthesizes a candidate protocol via the Evolution Engine. Crucially, before deployment, the candidate logic must pass a rigorous Immune System—a safety verification layer that performs formal checks and sandboxed validation. Upon approval, the system executes a millisecond-level Hot Swap on the underlying polymorphic containers. This architecture facilitates a fundamental paradigm shift from “rigid rules” to “organic adaptation,” ensuring that the network can emerge into optimal communication forms in response to environmental flux while maintaining strict endogenous security.
@@
The Darwin Cortex is the intelligent decision-making center where autonomous agents and LLMs reside [23, 28]. It functions as the “Slow Path” (System 2), responsible for high-level reasoning, intent sensing, and protocol synthesis [29]. By analyzing environmental pressure and user intent, Layer 2 generates the specific protocol “mutations” (bytecode) that are then pushed down to Layer 1 for execution [32, 16].
@@ IV-B The Evolutionary Loop
    Mutation (LLM Generation): When a performance bottleneck or intent mismatch is detected, the Darwin Cortex triggers a “directed mutation.” Unlike random biological mutations, the LLM utilizes its vast knowledge base to synthesize a specific protocol snippet—tailored to the current context—and compiles it into restricted WASM bytecode [5].
@@
    Hot Swap (Deployment): Once consensus is reached, the Fluid Cortex performs a millisecond-level logic reload. The old protocol stack is “melted” and replaced by the new, optimized bytecode, completing the evolutionary leap without packet loss or session termination.

### 13 Apr 2026 v2

@@ I Introduction
For decades, the classical edifice of computer science has been built upon the aesthetics of completeness and determinism. Nowhere is this more evident than in the narrow waist of the TCP/IP architecture. While this rigid design enabled the global expansion of the Internet, it has inadvertently led to a phenomenon known as protocol ossification [1, 9]. In modern heterogeneous environments—ranging from high-latency satellite constellations to massive-scale IoT deployments—the standardization cycle of traditional protocols is prohibitively long, often taking years to ratify minor adjustments to headers or congestion control logic. Furthermore, the traditional paradigm of Code is Law assumes that engineers can pre-emptively define every combination of input and state during the design phase. However, this reliance on mechanical completeness creates inherently brittle systems. When faced with undefined edge cases or novel environmental stressors, rule-based systems either crash or fall into unknown error states, lacking the flexibility to adapt without human intervention.
As we enter the era of Large Language Models (LLMs) and autonomous Agents, the fundamental nature of a network node is changing [27, 20]. Computing nodes are no longer merely passive executors of deterministic instructions; they have evolved into intelligent agents capable of probabilistic reasoning and intent understanding. In this context, legacy deterministic communication protocols have become a bottleneck, sacrificing efficiency for human readability—a constraint we term Carbon Hegemony [^1] Here Carbon Hegemony refers to the constraints imposed by human-centric cognitive structures on machine-to-machine communication, rather than environmental carbon emissions.. DarwinNet seeks to dismantle this by envisioning a formal parting of ways between computer protocols and human language, allowing machines to communicate via high-dimensional logic that transcends human cognitive structures. Crucially, this evolution does not imply the abandonment of existing network foundations. Instead, DarwinNet is positioned as a high-order augmentation for the 6G era, designed to handle hyper-heterogeneous edge scenarios where traditional manual standardization reaches its scalability limits, while still relying on the global reachability provided by the TCP/IP narrow waist as its foundational anchor.
@@
To realize the transition from structural fragility to evolutionary anti-fragility depicted in Fig. 1, this paper introduces DarwinNet—a bio-inspired, agent-driven network architecture that enables protocols to be grown rather than designed. Unlike traditional networks that succumb to ossification under static standards, DarwinNet conceptualizes the network as a living ecosystem capable of autonomous adaptation [24]. Structurally, this is achieved through a tri-layered decoupling: a Physical/Mathematical Anchor (Layer 0) for immutable constraints, a Fluid Cortex (Layer 1) serving as the adaptive execution body, and a Darwin Cortex (Layer 2) acting as the intelligent decision-making brain. By shifting the paradigm from rigid header-matching to high-dimensional intent negotiation, DarwinNet allows nodes to synthesize and deploy specialized communication logic in real-time. This mechanism ensures that network anomalies are no longer fatal errors, but rather catalysts for evolution that drive the system toward higher efficiency and survival.
In summary, DarwinNet represents a fundamental reconstruction of the network interaction model, moving from rigid rules to organic adaptation [11]. The core contributions of this work are as follows: 1. Paradigm Shift from Design to Growth:We propose a self-adaptive network architecture that achieves anti-fragility. By allowing protocols to evolve as fluid expressions of node intents, the network can autonomously adapt to environmental mutations—such as unforeseen congestion or novel attack patterns—without human-issued patches. 2. Intent-to-Bytecode (I2B) Synthesis with Dual-Loop Mechanism:We design an adaptive framework that leverages LLMs to transform high-level business intents into restricted, executable WebAssembly (WASM) bytecode. This is governed by a dual-loop architecture: an outer loop for cognitive protocol synthesis and an inner loop for millisecond-level Hot Swaps within a zero-trust sandbox [18]. 3. Protocol Solidification Index (PSI):We define a novel metric to quantify the maturity of the evolving system. This index captures the process of Protocol Liquefaction, where high-level reasoning (System 2) gradually collapses into highly efficient, compiled execution paths (System 1) as common interaction patterns stabilize. 4. Reliability Verification Paradigm:We introduce a mathematical verification framework based on the Crow-AMSAA (Duane) model. By treating protocol mutations as a reliability growth process, we provide a formal method to ensure that autonomous evolution remains within safety boundaries while continuously reducing failure rates over time.
@@ II Related Work
Programmable Data Planes and eBPF. The advent of P4 (Programming Protocol-independent Packet Processors) and eBPF (extended Berkeley Packet Filter) has shifted the networking paradigm from fixed-function hardware to programmable targets [4]. These technologies allow developers to define custom packet-processing logic and execute it at near-native speeds within the kernel or on specialized switching chips [4, 18]. However, existing programmable data planes primarily function as execution environments that require human-defined logic at compile-time or deployment-time. While they provide the necessary body for flexible networking, they lack an autonomous brain capable of real-time reasoning and logic synthesis [19]. DarwinNet builds upon these execution primitives but elevates them by introducing an agent-driven layer that dynamically generates and swaps protocol logic during runtime.
@@
Intent-Based Networking (IBN). Traditional IBN systems aim to bridge the gap between high-level business goals and low-level network configurations [15]. Most current IBN implementations are limited to parameter-level adjustments, such as automatically tuning TCP window sizes or re-routing paths based on pre-defined policies [19]. DarwinNet pushes the boundaries of IBN into the domain of logic-level synthesis [21, 20]. Instead of mapping intents to a library of static, human-written scripts, DarwinNet utilizes a fluid cortex to synthesize entirely new protocol behaviors in the form of executable bytecode. This transition from Intent-to-Policy to I2B enables a degree of architectural agility that conventional IBN cannot achieve [15, 5].
@@
Emergent Communication in Multi-Agent Reinforcement Learning (MARL). Research in MARL has demonstrated that autonomous agents can invent their own communication protocols to solve collaborative tasks, often resulting in high-dimensional vector-based languages that are unintelligible to humans [16, 22, 7, 17]. While these emergent protocols achieve high efficiency, they are often black boxes embedded within neural weights, making them difficult to verify, port, or debug. While maintaining safety via WASM, DarwinNet acknowledges that human experts are no longer the primary coders of protocol syntax. This marks a transition of computer science from a pure engineering discipline to an experimental science, where humans act as gardeners of an evolving ecosystem rather than micro-managers of bit-level headers. By using WASM as the synthesis target [31, 21], DarwinNet ensures that evolved protocols are explicit, portable, and subject to formal safety constraints within a zero-trust sandbox [31].
@@
Mimic Defense and Endogenous Security. The theory of Mimic Defense, characterized by Dynamic Heterogeneous Redundancy (DHR), suggests that system security can be enhanced by introducing uncertainty and dynamic transformations to thwart attackers. DarwinNet inherently realizes this security philosophy through the mechanism of protocol mutation [2]. By constantly evolving the syntax and semantics of communication—rather than relying on a static, standardized instruction manual like TCP/IP—DarwinNet creates a Moving Target Defense (MTD) at the protocol level. This dynamic nature significantly increases the cost of reconnaissance and exploitation for attackers, as the network’s attack surface is in a state of constant, intelligent flux.
@@
Recursive InterNetwork Architecture (RINA). RINA represents a fundamental rethink of the OSI model, proposing a recursive structure that treats networking as a single type of layer (the Distributed IPC Facility) repeated at different scales [12, 10]. While RINA provides an elegant theoretical framework for cross-layer optimization and de-layering, it still assumes a relatively stable set of mechanisms and policies. DarwinNet shares RINA’s vision of breaking down rigid layer boundaries but focuses more on the biological growth of functions rather than the structural recursion of the architecture [30]. DarwinNet’s liquefaction process treats protocol functions as atomic components that can be organically reassembled and evolved based on environmental fitness rather than predefined recursive templates.
@@ III-A Dual-Process Theory in Network Intelligence
The evolution of a protocol in DarwinNet is essentially the process of liquid System 2 reasoning collapsing into solid System 1 reflexes to achieve maximum efficiency [29, 23, 3].
@@ III-B WebAssembly and Sandbox Isolation
To ensure that autonomously synthesized logic does not compromise node security, we utilize WASM as the fluid execution target. WASM provides a memory-safe, hardware-independent, and near-native performance environment [18]. Its capability for millisecond-level instantiation and hot-swapping allows DarwinNet to perform logic mutations without interrupting active data streams. Furthermore, the use of restricted WASM bytecode facilitates formal verification of synthesized logic, ensuring it adheres to the Constitution defined in Layer 0 [31].
@@ III-C Crow-AMSAA Reliability Growth Model
The evolutionary maturity of DarwinNet is modeled using the Crow-AMSAA (Duane) Power Law process. This model describes the reliability growth of a system undergoing iterative discovery and repair cycles [8]. The failure (or protocol mismatch) intensity function is defined as:
@@ IV The Proposed Framework
This tri-layered, dual-path design ensures that DarwinNet can decouple the intelligence of the protocol from the mechanics of the transmission, allowing nodes to evolve their communication language without disrupting the underlying physical connectivity.
@@ IV-A Layered Architecture: From Anchor to Fluid
DarwinNet is organized into three distinct layers, transitioning from Physical/Mathematical Hardness at the base to Intelligent Liquidity at the apex.
@@
Layer 0 represents the fundamental laws of physics of the network that remain constant throughout the evolutionary process. This layer ensures basic bit-level connectivity (e.g., the standard TCP/IP stack for physical reachability) and enforces mathematical truths, such as cryptographic verification and logical consistency ($1+1=2$). In this paradigm, we relinquish micro-control over the communication process (the Silicon-driven protocol) and retain only mathematical constraints over the results (the Consensus). Layer 0 ensures that mathematics remains the only universal language between carbon-based intent and silicon-based execution, serving as the final tether of human control over autonomous evolution [10].
This layer functions as the Mathematical Constitution of the network. No matter how complex or liquid the evolved logic in Layer 1 becomes, it is strictly forbidden from violating the physical and security predicates pre-defined by human architects in Layer 0. This ensures that the autonomy of the network is always bounded by a Human-in-the-loop safety framework, preventing any unpredictable or runaway protocol behavior.
@@
Layer 1 is the execution environment of the network, characterized by its liquid state. Unlike traditional fixed-protocol implementations, the Fluid Cortex utilizes a high-performance, zero-trust sandbox (e.g., WebAssembly) to run protocol logic. This layer serves as the system’s body or the Fast Path (System 1), executing synthesized bytecode at near-native speeds. It supports millisecond-level hot-swapping of logic—such as switching from a JSON-based parser to a specialized binary delta-compression algorithm—without interrupting data flows.
@@
As illustrated in Fig. 3, the Fluid Cortex is architected as a bionic Dual-Loop feedback system, conceptualizing the network node as a self-evolving liquid organism [25, 30]. This design explicitly bifurcates the processing logic into two cognitive pathways:
  • The Polymorphic Runtime (System 1): Occupying the lower tier, this body serves as the Fast Path. It utilizes deformable, liquid-like containers to carry real-time data streams. Its primary objective is the pursuit of ultimate execution efficiency, operating on the evolved reflexes (bytecode) generated by the upper layer.
  • The Darwin Cortex (System 2): Occupying the upper tier, this brain serves as the Slow Path. It is responsible for high-level perception, identifying shifts in user intent and environmental anomalies through its sensory nerves.
The synergy between these two loops enables a sophisticated evolutionary reflex: when the sensory nerves trigger an evolution request, the Darwin Cortex retrieves historical experiences from a Gene Pool and synthesizes a candidate protocol via the Evolution Engine. Crucially, before deployment, the candidate logic must pass a rigorous Immune System—a safety verification layer that performs formal checks and sandboxed validation. Upon approval, the system executes a millisecond-level Hot Swap on the underlying polymorphic containers. This architecture facilitates a fundamental paradigm shift from rigid rules to organic adaptation, ensuring that the network can emerge into optimal communication forms in response to environmental flux while maintaining strict endogenous security.
@@
The Darwin Cortex is the intelligent decision-making center where autonomous agents and LLMs reside [23, 28]. It functions as the Slow Path (System 2), responsible for high-level reasoning, intent sensing, and protocol synthesis [29]. By analyzing environmental pressure and user intent, Layer 2 generates the specific protocol mutations (bytecode) that are then pushed down to Layer 1 for execution [32, 16].
@@ IV-B The Evolutionary Loop
    Mutation (LLM Generation): When a performance bottleneck or intent mismatch is detected, the Darwin Cortex triggers a directed mutation. Unlike random biological mutations, the LLM utilizes its vast knowledge base to synthesize a specific protocol snippet—tailored to the current context—and compiles it into restricted WASM bytecode [5].
@@
    Hot Swap (Deployment): Once consensus is reached, the Fluid Cortex performs a millisecond-level logic reload. The old protocol stack is melted and replaced by the new, optimized bytecode, completing the evolutionary leap without packet loss or session termination.